# Self-regularized entropy: How much do quantum black holes violate the no-hair theorem?

### v1

@@ Abstract
We compute the canonical (brick wall) entropy of Hawking radiation in a quantum black hole whose exterior is described, to first order in a small quadrupole parameter, by the static $q$-metric, which is an exact vacuum solution of the Einstein equations. Counting near horizon quasinormal modes shows that a modest quadrupolar deformation self-regularizes the ultraviolet divergence: the entropy of Hawking radiation is finite for any non-vanishing quadrupole, without an ad hoc cutoff. Matching this canonical entropy to the Bekenstein-Hawking entropy leads to no-hair violating multipoles, at percent-to-tens-of-percent level, and provides concrete observational targets for the Next Generation Event Horizon Telescope (ngEHT) and the Laser Interferometer Space Antenna (LISA).
@@ I Introduction
Despite the success of black hole thermodynamics, the microscopic origin of the Bekenstein-Hawking entropy remains unsettled; a variety of horizon degrees of freedom have been proposed e.g., PhysRevD.7.2333 ; 1998PhRvL..81.4293H ; 1996PhLB..379…99S ; 1997NuPhS..57…65I ; 2000gr.qc…..5126A ; 2003PhRvL..90h1301D ; 2004CQGra..21.5245M ; 2006CQGra..23.3849D ; Saravani:2012is ; 2021PhRvD.104l1502C ; 2023PhLB..84137901O ; 2023arXiv230404491N , yet no single picture is universally accepted. In this work, we explore a conservative, geometry-based option: we treat the quasinormal (QN) modes as an effective set of near horizon excitations that can account for (at least part of) the entropy. The idea is to model quantum fluctuations of the spacetime geometry as superpositions of QN mode eigenstates, in close analogy with the quantization of a harmonic oscillator, where the occupation of each mode provides a natural bookkeeping of microstates.
@@
The brick wall model of ’t Hooft tHooft:1984kcu ; tHooft:1996rdg remains the simplest way to estimate the canonical (thermodynamic) entropy of quantum fields around a black hole, but in the exactly spherical Schwarzschild geometry the mode sum diverges logarithmically and one is forced to introduce an ad‑hoc proper distance cutoff. In this paper, we show that the brick wall ultraviolet divergence disappears once the Schwarzschild exterior is perturbed by an arbitrarily small quadrupolar deformation.
@@
An implication of a brick wall model for a quantum black hole is that we can send and receive signals from the quantum structure near the horizon, within a finite time. This has motivated searches for “echoes” following black hole mergers in gravitational wave e.g, Abedi:2016hgu ; Cardoso:2016rao ; Oshita:2019sat ; Wang:2019rcf ; Oshita:2023tlm ; Vellucci:2022hpl ; Arrechea:2024nlp (see Abedi:2020ujo ; Berti:2025hly for reviews). However, this would also imply that a stationary spacetime may not obey the no-hair theorem, as it can learn about the absence of a regular horizon in a finite time (e.g., Bah:2021jno ). In this paper, we propose that the $q$-metric may be used as a way to probe the impact of these near horizon quantum effects on the rest of the exterior spacetime, through solving the Einstein field equations.
@@
Our goal is to isolate the single geometric ingredient that makes the Schwarzschild brick wall sum diverge and to test how a minimal, physically motivated deformation changes the near horizon density of states. The static Zipoy-Voorhees $q$-metric osti_4201189 ; PhysRevD.39.2904 is ideal for this. It is an exact, asymptotically flat, one parameter deformation of Schwarzschild that remains a vacuum solution of Einstein’s equations and reduces smoothly to Schwarzschild as $q\,\to\,0$. The single parameter $q$ encodes a quadrupolar departure from spherical symmetry; the first nontrivial hair allowed once perfect symmetry is relaxed. Working in vacuum general relativity avoids modeling ambiguities from matter sources or modified field equations and keeps all changes traceable to geometry alone.
We work with the $q$-metric line element e.g., osti_4201189 ; PhysRevD.39.2904 , a one parameter ($q$) vacuum deformation of Schwarzschild, and restrict to first order in $|q|\ll 1$. The transverse circumference pinches (for $q>0$) and the density of states softens. A direct count of quasinormal modes shows that the canonical entropy converges for a range of $q>0$; no brick wall cutoff or stretched horizon dissipation is required. Thus, the $q$-metric delivers a closed form, analytically tractable test of self-regularisation without introducing ad hoc cutoffs or dissipation. Treating $q$ as the coarse grained imprint of horizon microstates and imposing that the entropy radiated by Hawking quanta never exceeds the Bekenstein-Hawking entropy fixes its natural scale to $|q|\simeq 0.1$ for all astrophysical black holes. A distortion of this magnitude is not merely theoretical: it shifts the lowest mass and current multipole moments linearly in $q$, supplying a concrete observational target, e.g., for next‑generation horizon imaging (ngEHT) and gravitational wave spectroscopy (LISA). We also show that in the stationary extension of the $q$-metric the same parameter $q$ induces linear, Kerr-violating shifts in the lowest multipoles captured by two Kerr null combinations $Q_{1}$ and $Q_{2}$, so that the entropy motivated bound on $|q|$ becomes an observational target for ngEHT and LISA.
We emphasise that we are not claiming astrophysical objects are exact $q$-metrics. Rather, the $q$-metric is the simplest vacuum laboratory that (a) preserves the universal Rindler $\times$ transverse near horizon structure, (b) introduces precisely one controlled deformation that governs the ultraviolet behaviour of the mode sum, and (c) connects directly to measurable multipoles. More general ”bumpy” metrics or rotating spacetimes introduce additional scales (e.g. superradiance, spin couplings) that obscure this single parameter test without changing the qualitative conclusion: any axisymmetric deformation, self-regularises the brick wall in the same way; the $q$-metric is simply the cleanest analytic example.
@@
The rest of the manuscript is orgnaized as follows: Section II reviews the $q$-metric. In Section III we derive the scalar wave operator and the exact radial Klein-Gordon problem, while Section IV performs the WKB counting and presents the entropy. Section V offers observational prospects for measuring quadrupole, and Section VI gathers the main implications. In this work, we adopt the convention $G=c=\hbar=k_{B}=1$.
II Spacetime
@@
The $q$-metric describes static, axially symmetric, and asymptotically flat exact solutions to the vacuum Einstein field equations. By incorporating the quadrupole moment, the $q$-metric accounts for the deformation of the gravitational field which is not captured by the spherically symmetric Schwarzschild solution. Consequently, the $q$-metric characterizes the exterior gravitational field of an isolated static axisymmetric object. The presence of a quadrupole can change the geometric properties of spacetime and affect the timelike and lightlike trajectories e.g., 2021A&A…654A.100F ; 2025PhRvD.111d5003A ; 2025A&A…699A.266F ; 2025EPJC…85..148F . As a result, the $q$-metric is a particularly useful toy model for the gravitational fields of realistic astrophysical bodies where the effects of higher-order multipole moments cannot be ignored. The metric in Schwarzschild-like coordinates $(t,x,y,\phi)$ is presented as follows osti_4201189 ; PhysRevD.39.2904
@@
where $t\in(-\infty,+\infty)$, $r\in(2m,+\infty)$, $\theta\in[0,\pi]$, and $\phi\in[0,2\pi)$, also $m$ is a parameter which is expressed in the dimension of length. This metric contains two free parameters, namely $m$, and quadrupole parameter $q$. In the case of vanishing $q$, the Schwarzschild metric is recovered. Furthermore, values of $q>0$ describe the exterior field of an oblate central object, while $q<0$ is related to a prolate source. Additionally, the $q$-metric, aside from a central curvature singularity at $r=0$, has another singularity located at a finite distance from the origin at $r=2m$ for any chosen nonzero value of quadrupole. By Geroch’s definition (1970JMP….11.2580G, ), for multipole moment and for avoiding a negative mass distribution, the quadrupole parameter is restricted to the domain of $q\in(-1,\infty)$.
@@ II.1 Echo time and entropy bound
$\simeq 8mq^{-2}+{\cal O}(mq^{-1})=\frac{1}{\pi T_{H}q^{2}}.$
We take the asymptotic Hawking temperature to be that of a Schwarzschild black hole with the same mass, $T_{H}=1/(8\pi m)$. This is a modeling choice consistent with the small-$q$ limit and with our use of the near horizon cavity to count modes Oshita_2023 . The $\theta$ dependence in the integrand produces only an $\mathcal{O}(m)$ modulation and does not affect the leading $mq^{-2}$ scaling  [^1] To first order in $q$, the light-ring radius shifts to $r_{\rm LR}=m(3+2q)$; using $3m$ as the upper limit changes $t_{\rm echo}$ only at $\mathcal{O}(q^{0})$, hence subleading to the $q^{-2}$ term.. We note that for the entropy in the Hawking radiation not to exceed the Bekenstein-Hawking horizon entropy of the black hole Oshita_2023 , we require
$t_{\rm echo}\lesssim\frac{\ln S}{2\pi T_{H}}\Rightarrow q^{2}\gtrsim\frac{2}{\ln S}.$ (3)
This implies that, while any value of $q$ solves the classical vacuum Einstein equations, the finite entropy of quantum effects that lead to Hawking radiation puts a lower bound on $|q|$, that would describe a semi-classical spacetime. This sets the minimal modifications we expect from Schwarzschild/Kerr solutions, even far from the horizon.
Equation (3) should be read as a conservative lower bound on $q$. It is obtained by treating the Hawking flux and surface gravity as those of Schwarzschild (i.e. $T_{H}=1/8\pi M$) and by using the leading small $|q|$ echo time scaling, so that the total entropy trapped within the angular momentum barrier does not exceed the Bekenstein-Hawking value Oshita_2023 . This estimate merely sets the minimal distortion required to avoid overcounting. In what follows (Sections III-V), we compute the canonical entropy directly in the $q$-metric by near horizon WKB mode counting, which yields a sharper relation $q=q(S)$ for this spacetime and thus a more accurate bound or estimate on $q$.
@@ III.1 Separation of variables
Because the angular operator carries unequal $\rho^{\mp 2q}$ weights in the $\theta$ and $\phi$ parts, exact separation into a single ODE holds rigorously in the axisymmetric sector $m=0$. However, for general $m$ and $|q|\ll 1$ one may use the $\ell$-dominated approximation. Replacing the weighted angular operator by its $\ell(\ell+1)$ eigenvalue; this is accurate at leading order in $q$ (the corrections are $O(q\,/\ell)$ and do not affect the near horizon exponents or the WKB scaling).
Projecting (10) onto $Y_{\ell m}$ yields the radial equation
@@
Using $\rho=\left(\frac{\beta\,r_{*}}{(4m)^{\alpha/2}}\right)^{\!1/\beta}$ and noting $\alpha-2=-2\beta$, the curvature piece becomes a universal inverse square wall,
@@ III.2 Classical turning points and cavity length
The WKB approximation is valid in the domain where the radial wave number $k_{r}^{2}\,=\,\omega^{2}-V_{\ell}(\rho)$ is positive, i.e. between two classical turning points. For the potential $V_{\ell}(\rho)$ in equation (17) the inner edge is $\rho_{\min}=0$, while the outer edge $\rho_{\max}$ is defined by $k_{r}^{2}(\rho_{\max})=0$:
@@
(valid for  $\ell\,\gg\,1$). Setting $\omega^{2}=f(\rho_{\max})\,k_{\perp}^{2}(\rho_{\max})$ gives
@@
Cavity length.
Since the inner edge is at $r^{*}_{\min}=0$, the proper WKB cavity length is simply
$L(\omega,\ell)=r^{*}_{\max}.$ (26)
@@
Besides the centrifugal term, the effective potential contains the “curvature” part
$V_{\mathrm{curv}}(\rho)=f(\rho)\,\frac{\alpha\beta}{2\rho^{2}}=\frac{\alpha\beta}{32\,m^{2}}\left(\frac{\rho}{4m}\right)^{\alpha-2},\quad 0<\alpha<2,\ \beta>0,$ (27)
where we used $f(\rho)=\bigl{(}\rho/(4m)\bigr{)}^{\alpha}$ and $\alpha-2=-2\beta$. Since $\alpha<2$ for every $|q|<1$, one has $V_{\mathrm{curv}}\to+\infty$ as $\rho\to 0$, so the inner (classical) turning point is determined by
$\omega^{2}=V_{\mathrm{curv}}(\rho_{\mathrm{in}})$ $\;\;\Longrightarrow\;\;\left(\frac{\rho_{\mathrm{in}}}{4m}\right)^{\alpha-2}=\frac{32\,m^{2}\,\omega^{2}}{\alpha\beta}$
@@
$r^{*}_{\mathrm{min}}=\frac{4m}{\beta}\left(\frac{\rho_{\mathrm{in}}}{4m}\right)^{\beta}=\frac{4m}{\beta}\left(\frac{\alpha\beta}{32\,m^{2}\,\omega^{2}}\right)^{\!1/2}=\frac{1}{\omega}\,\sqrt{\frac{\alpha}{2\beta}}\approx\frac{1}{\omega q}\,$,
which tends to zero for large frequencies. Furthermore, $r^{*}_{\mathrm{min}}\ll r^{*}_{\mathrm{max}}$ for $q\ll 1$ and $\omega\sim T_{\rm H}$. Thus the divergent curvature term acts effectively as a hard wall at $r^{*}\approx 0$. Replacing $\rho_{\mathrm{in}}$ by 0 therefore shifts the cavity length $L$ only by an $\mathcal{O}(q)$ fraction and does not affect the WKB scaling or the finiteness of the entropy. This justifies setting $\rho_{\min}=0$ in equation (26).
For $q=0$ (Schwarzschild) equation (25) reduces to the familiar logarithmic behavior $L\sim 4m\ln\big(\ell/(m\omega)\big{missing})$+ const. For any $|q|\ll 1$, it becomes a power law in $\omega/\ell$ displayed by equation (25); this change of scaling underlies the self-regularisation of the mode sum discussed in the next section.

### v2

# Self-regularized entropy: What does black hole entropy predict for tests of Kerr no-hair theorem?

@@ Abstract
We compute the canonical (brick-wall) entropy of Hawking radiation in a in a quantum black hole model whose strong-field exterior is modeled phenomenologically, to first order in a small quadrupole parameter, by the static q-metric, which is an exact vacuum solution of the Einstein equations. WKB counting of trapped near-horizon cavity modes shows that, within the perturbative small-deformation regime studied here, a modest quadrupolar deformation self-regularizes the ultraviolet divergence: the entropy becomes finite without an ad hoc cutoff. Adopting the Hawking temperature and the Bekenstein-Hawking entropy of a Schwarzschild black hole of the same mass as external thermodynamic inputs, matching this canonical entropy to that benchmark yields an entropy-motivated deformation scale which, when interpreted phenomenologically in a stationary extension, corresponds to percent-to-tens-of-percent violations of the Kerr multipole relations, and provides concrete observational targets for the Next Generation Event Horizon Telescope (ngEHT), the Laser Interferometer Space Antenna (LISA), and planned third-generation (3G) ground-based gravitational wave observatories.
@@ I Introduction
Despite the success of black hole thermodynamics, the microscopic origin of the Bekenstein-Hawking entropy remains unsettled; a variety of horizon degrees of freedom have been proposed e.g., Bekenstein (1973); Hod (1998); Strominger and Vafa (1996); Immirzi (1997); Ashtekar et al. (2000); Dreyer (2003); Meissner (2004); Dittrich and Loll (2006); Saravani et al. (2015); Cadoni et al. (2021); Oshita and Afshordi (2023a); Nam (2023), yet no single picture is universally accepted. In this work, we explore a conservative, geometry-based option: we ask whether a minimal deformation of the near-horizon geometry can render the brick-wall mode count finite, thereby providing a controlled arena for horizon scale microphysics. Concretely, we focus on the high frequency near-horizon cavity (trapped) modes that dominate the brick-wall entropy in echo capable, horizonless compact objects, and we distinguish them from classical black hole quasinormal modes defined by purely ingoing conditions at the horizon and outgoing behaviour at infinity.
@@
The brick-wall model of ’t Hooft ’t Hooft (1985, 1996) remains the simplest way to estimate the canonical (thermodynamic) entropy of quantum fields around a black hole, but in the exactly spherical Schwarzschild geometry the mode sum diverges logarithmically and one is forced to introduce an ad‑hoc proper distance cutoff. In this paper, we show that the brick-wall ultraviolet divergence is self-regularized once the Schwarzschild exterior is perturbed by a sufficiently small but nonzero quadrupolar deformation. In this paper, we show that the brick-wall ultraviolet divergence disappears once the Schwarzschild exterior is perturbed by an arbitrarily small quadrupolar deformation.
@@
An implication of a brick-wall model for a quantum black hole is that we can send and receive signals from the quantum structure near the horizon, within a finite time. This has motivated searches for “echoes” following black hole mergers in gravitational wave e.g, Abedi et al. (2017); Cardoso et al. (2016); Oshita et al. (2020b); Wang et al. (2020); Oshita and Afshordi (2023b); Vellucci et al. (2023); Arrechea et al. (2024) (see Abedi et al. (2020); Berti and others (2025) for reviews). However, this would also imply that a stationary spacetime may not obey the no-hair theorem, as it can learn about the absence of a regular horizon in a finite time (e.g., Bah et al. (2021)). In this paper, we use the $q$-metric as an analytic vacuum laboratory for this purpose: it introduces a single controlled quadrupolar deformation while keeping the exterior governed by the Einstein equations.
@@
Our goal is to isolate the single geometric ingredient that makes the Schwarzschild brick-wall sum diverge and to test how a minimal, physically motivated deformation changes the near horizon density of states. The static Zipoy-Voorhees $q$-metric Erez and Rosen (1959); Quevedo (1989) is ideal for this. It is an exact, asymptotically flat, one parameter deformation of Schwarzschild that remains a vacuum solution of Einstein’s equations and reduces smoothly to Schwarzschild as $q\,\to\,0$. The single parameter $q$ encodes a quadrupolar departure from spherical symmetry; the first nontrivial hair allowed once perfect symmetry is relaxed. Working in vacuum general relativity avoids modeling ambiguities from matter sources or modified field equations and keeps all changes traceable to geometry alone.
We work with the static $q$-metric line element Erez and Rosen (1959); Quevedo (1989), a one-parameter vacuum deformation of Schwarzschild, and restrict to first order in $|q|\ll 1$. Using the brick-wall framework, we perform a WKB count of the high-frequency trapped modes in the near-horizon cavity and compute the associated canonical entropy in closed form. We find that the Schwarzschild logarithmic divergence is removed: for sufficiently small nonzero $|q|$ in the regime considered here, the mode sum converges without introducing an ad hoc proper-distance cutoff. Treating the Hawking temperature and the Bekenstein–Hawking entropy of a Schwarzschild black hole of the same mass as external thermodynamic benchmarks, we obtain an entropy-motivated deformation scale of order $|q|\sim 0.2$ for astrophysical black holes. In a stationary extension of the $q$-metric, the same parameter induces Kerr-violating shifts in the lowest multipoles (captured by the Kerr-null combinations $Q_{1}$ and $Q_{2}$), making this scale an observational target for ngEHT and LISA.
For nonzero $q$, the static $q$-metric is horizonless and contains a singularity at the would-be horizon, so it is not interpreted here as a literal classical black-hole exterior. Instead, it is used only as an effective description of the strong-field exterior region of a quantum black hole model, with the microscopic completion of the near-horizon region left unspecified. In particular, the asymptotic Hawking temperature and the Bekenstein-Hawking entropy for a Schwarzschild black hole of the same mass are treated as external thermodynamic inputs. Within this framework, the role of the $q$-metric is to modify the near-horizon redshift and transverse geometry in order to regulate the WKB density of states responsible for thermodynamic entropy, in a manner consistent with the Einstein field equations. We do not claim that astrophysical objects are exact $q$-metrics. Rather, the $q$-metric provides the simplest vacuum setting that isolates a single quadrupolar deformation and allows one to track its consequences analytically; (a) preserves the approximate universal Rindler $\times$ transverse near horizon structure, (b) introduces precisely one controlled deformation that governs the ultraviolet behaviour of the mode sum, and (c) connects directly to observable multipoles. More general bumpy or rotating spacetimes introduce additional scales that may change quantitative details. However, they are not expected to alter the qualitative self-regularization mechanism introduced here and are thus left for future work.
@@
The rest of the manuscript is organized as follows: Section II reviews the $q$-metric. In Section III we derive the scalar wave operator and the exact radial Klein-Gordon problem, while Section IV performs the WKB counting and presents the entropy. Section V offers observational prospects for measuring quadrupole, and Section VI gathers the main implications. In this work, we adopt the convention $G=c=\hbar=k_{B}=1$.
II Space-time
@@
The $q$-metric describes static, axially symmetric, and asymptotically flat exact solutions to the vacuum Einstein field equations. By incorporating the quadrupole moment, the $q$-metric accounts for the deformation of the gravitational field which is not captured by the spherically symmetric Schwarzschild solution. Consequently, the $q$-metric characterizes the exterior gravitational field of an isolated static axisymmetric object. The presence of a quadrupole can change the geometric properties of spacetime and affect the timelike and lightlike trajectories e.g., Faraji and Trova (2021); Amaro and Faraji (2025); Faraji and Rosa (2025); Faraji (2025). As a result, the $q$-metric is a particularly useful toy model for the gravitational fields of realistic astrophysical bodies where the effects of higher-order multipole moments cannot be ignored. The metric in Schwarzschild-like coordinates $(t,x,y,\phi)$ is presented as follows Erez and Rosen (1959); Quevedo (1989)
@@
where $t\in(-\infty,+\infty)$, $r\in(2m,+\infty)$, $\theta\in[0,\pi]$, and $\phi\in[0,2\pi)$, also $m$ is a parameter which is expressed in the dimension of length. This metric contains two free parameters, namely $m$, and quadrupole parameter $q$. In the case of vanishing $q$, the Schwarzschild metric is recovered. Furthermore, values of $q>0$ describe the exterior field of an oblate central object, while $q<0$ is related to a prolate source. Additionally, the $q$-metric, aside from a central curvature singularity at $r=0$, has another singularity located at a finite distance from the origin at $r=2m$ for any chosen nonzero value of quadrupole. By Geroch’s definition (Geroch, 1970), for multipole moment and for avoiding a negative mass distribution, the quadrupole parameter is restricted to the domain of $q\in(-1,\infty)$.
@@ II.1 Echo time and entropy bound
$\simeq 8mq^{-2}+{\cal O}(mq^{-1})=\frac{1}{\pi T_{\rm H}q^{2}}.$
The $\theta$-dependence in the integrand produces only an $\mathcal{O}(m)$ modulation and does not affect the leading $mq^{-2}$ scaling  [^1] To first order in $q$, the light-ring radius shifts to $r_{\rm LR}=m(3+2q)$; using $3m$ as the upper limit changes $t_{\rm echo}$ only at $\mathcal{O}(q^{0})$, hence subleading to the $q^{-2}$ term.. We take the asymptotic Hawking temperature to be that of a Schwarzschild black hole with the same mass, $T_{\rm H}\simeq 1/(8\pi m)$, as an external thermodynamic input.
Motivated by the conjectured correspondence between the fast scrambling Sekino and Susskind (2008) and echo time scales Saraswat and Afshordi (2020); Oshita and Afshordi (2023b), we adopt the following conservative thermodynamic consistency hypothesis: the entropy associated with the trapped near-horizon thermal atmosphere should not parametrically exceed the Bekenstein-Hawking entropy of a Schwarzschild black hole of the same mass. Following Oshita and Afshordi (2023b), the echo delay time of a finite-entropy compact object with approximately Rindler near-surface geometry, admits an upper bound (the “quantum limit of silence”) of the form $t_{\rm echo}\lesssim\ln S/(2\pi T_{\rm H})$. Applying this bound to the $q$-metric echo time (2) yields
$q^{2}\gtrsim\frac{2}{\ln S}.$ (3)
Equation (3) should be read as a conservative rough estimate for a minimal $|q|$, required to avoid overcounting entropy.
In what follows (Sections III-V), we compute the canonical entropy directly in the $q$-metric by near-horizon WKB mode-counting, which yields a sharper relation $q=q(S)$ for this spacetime and thus a more accurate bound or estimate on $q$.
@@ III.1 Separation of variables
Because the angular operator carries unequal $\rho^{\mp 2q}$ weights in the $\theta$ and $\phi$ parts, exact separation into a single ODE holds rigorously in the axisymmetric sector $m=0$. For general $m$ we proceed by projecting onto the standard $Y_{\ell m}$ basis and using a large-$\ell$ (high angular momentum) approximation appropriate to the ultraviolet mode count. In this regime the detailed angular eigenfunctions can change, but the leading high-eigenvalue density is controlled by the transverse area (Weyl asymptotics) and is therefore insensitive to $O(1)$ anisotropies. Accordingly, we replace the weighted angular operator by an effective $\ell(\ell+1)$ eigenvalue at the level needed to capture the near-horizon WKB scaling; this may affect overall numerical prefactors but does not affect the power-law exponents that control convergence of the brick-wall sum. [^2] For each fixed $\rho$, the induced two-metric from (II.2) has area element $4m^{2}\sin\theta\,d\theta\,d\phi$, independent of $\rho$ and $q$ (since $g_{\theta\theta}g_{\phi\phi}=(4m^{2})^{2}\sin^{2}\theta$). By Weyl’s law, the asymptotic counting of large angular eigenvalues depends only on this area, so the ultraviolet scaling of the angular density of states is unchanged up to $O(1)$ factors.
Projecting (10) onto $Y_{\ell m}$ and using this approximation yields the radial equation
@@
Using $\rho=\left(\frac{\beta\,r_{*}}{(4m)^{\alpha/2}}\right)^{\,1/\beta}$ and noting $\alpha-2=-2\beta$, the $f$-derivative (Liouville) term becomes a universal inverse-square wall,
@@
It is worth emphasizing that the nature and boundary conditions used to define QNMs in exotic compact objects (ECO), radically differ from those of classical black holes, as such objects lack regular horizons (e.g., Cardoso et al. (2016); Wang et al. (2020)). Unlike the black hole QNMs that are mainly defined by the centrifugal barrier, the ECO QNMs are generically trapped near the would-be horizon, due to an inner Dirichlet-type boundary condition. As such, we focus on the near-horizon cavity: the region between the inner brick-wall/singularity (modeling the onset of quantum microstructure) and the outer centrifugal barrier. In this finite interval, the modes behave as standing waves, and the WKB condition is the usual Bohr-Sommerfeld quantisation for a bounded Schrödinger problem. The resulting mode density agrees with the high-frequency / large-$\ell$ QNM density up to sub-leading $\mathcal{O}(1)$ shifts in the overtone numbers, which do not affect the ultraviolet behaviour of the brick-wall entropy.
It is worth emphasizing that the boundary conditions relevant for horizonless exotic compact objects (ECOs) differ from those of classical black holes (e.g., Cardoso et al. (2016); Wang et al. (2020)). Classical black-hole quasinormal modes are defined by purely ingoing behaviour at the horizon and outgoing behaviour at infinity, whereas an ECO with an inner reflective (Dirichlet-type) boundary supports a near-horizon cavity whose resonant modes are trapped between the inner boundary (represented by the inner WKB turning point at the end of the near-horizon region, modeling the onset of quantum microstructure) and the outer centrifugal barrier. For our purposes we focus on the high-frequency standing-wave spectrum in this finite cavity, for which the WKB condition is the usual Bohr-Sommerfeld quantisation of a bounded Schrödinger problem. This approximation captures the leading large-$\ell$ / high-frequency mode density relevant for the ultraviolet behaviour of the brick-wall entropy; subleading $O(1)$ phase shifts and the small imaginary parts associated with leakage through the barrier do not affect the convergence properties of the mode sum.
@@ III.2 Classical turning points and cavity length
The WKB approximation is valid in the domain where the radial wave number $k_{r}^{2}\,=\,\omega^{2}-V_{\ell}(\rho)$ is positive, i.e. between two classical turning points. For the potential $V_{\ell}(\rho)$ in equation (17), the outer turning point $\rho_{\max}$ is defined by $k_{r}^{2}(\rho_{\max})=0$. The inner turning point $\rho_{\mathrm{in}}(\omega)$ is set by the inverse-square term in the effective potential (discussed below) and approaches $\rho=0$ in the high-frequency regime relevant for the ultraviolet mode count.
@@
Setting $\omega^{2}=f(\rho_{\max})\,k_{\perp}^{2}(\rho_{\max})$; hence for $\ell\gg 1$ the inverse-square term in (17) is subleading at the outer turning point and is neglected in determining $\rho_{\max}$. We obtain
@@
Besides the centrifugal term, the effective potential contains the inverse-square contribution generated by the $f$-derivative terms in equation (15),
$V_{\mathrm{L}}(\rho)=f(\rho)\,\frac{\alpha\beta}{2\rho^{2}}=\frac{\alpha\beta}{32\,m^{2}}\left(\frac{\rho}{4m}\right)^{\alpha-2},\quad 0<\alpha<2,\ \beta>0,$ (26)
where we used $f(\rho)=\bigl(\rho/(4m)\bigr)^{\alpha}$ and $\alpha-2=-2\beta$. Since $\alpha<2$ for every $|q|<1$, one has $V_{\mathrm{L}}\to+\infty$ as $\rho\to 0$, so the inner (classical) turning point is determined by
$\omega^{2}=V_{\mathrm{L}}(\rho_{\mathrm{in}})$ $\;\;\Longrightarrow\;\;\left(\frac{\rho_{\mathrm{in}}}{4m}\right)^{\alpha-2}=\frac{32\,m^{2}\,\omega^{2}}{\alpha\beta}$
@@
$r^{*}_{\mathrm{min}}=\frac{4m}{\beta}\left(\frac{\rho_{\mathrm{in}}}{4m}\right)^{\beta}=\frac{4m}{\beta}\left(\frac{\alpha\beta}{32\,m^{2}\,\omega^{2}}\right)^{\,1/2}=\frac{1}{\omega}\,\sqrt{\frac{\alpha}{2\beta}}\approx\frac{1}{\omega|q|}\,$,
which tends to zero for large frequencies. Since $\omega\,r^{*}_{\mathrm{min}}=\sqrt{\alpha/(2\beta)}$ is independent of $\omega$, keeping a nonzero $r^{*}_{\mathrm{min}}$ only shifts the Bohr-Sommerfeld overtone number by an $\omega$-independent constant and therefore does not affect the WKB density of states $\partial n/\partial\omega$, the scaling with $\omega/\ell$, or the ultraviolet convergence of the brick-wall entropy. Accordingly, we may set $\rho_{\min}=0$ (equivalently $r^{*}_{\min}=0$) in (28) at the level relevant for the ultraviolet mode count.
Cavity length.
The proper WKB cavity length is
$L(\omega,\ell)=r^{*}_{\max}-r^{*}_{\min}.$ (28)
We just obtained $r^{*}_{\min}\to 0$ at high frequencies and, moreover, it only induces an $\omega$-independent shift in the overtone number, so for the ultraviolet scaling we may set $L(\omega,\ell)\simeq r^{*}_{\max}$.
[…]